how much water (0% strength) should be added to 8L of 6% saline solution to reduce the strength to 4%?
How do I solve problems like this one?

Answer by stanbon(75887)

how much water (0% strength) should be added to 8L of 6% saline solution to reduce the strength to 4%?
Equation:
salt + salt = salt
0.06*8 + 0*x = 0.04(8+x)
--------
6*8 + 0 = 4*8 + 4x
-----
2*8 = 4x
x = 4 L (amt. of water to add)
====================================
Cheers,
Stan H.
================